# Scalable Temporal Motif Densest Subnetwork Discovery

### v1

@@ Abstract.
In this work we introduce the novel problem of identifying the temporal motif densest subnetwork, i.e., the densest subnetwork with respect to temporal motifs, which are high-order patterns characterizing temporal networks. This problem significantly differs from analogous formulations for dense temporal (or static) subnetworks as these do not account for temporal motifs. Identifying temporal motifs is an extremely challenging task, and thus, efficient methods are required. To this end, we design two novel randomized approximation algorithms with rigorous probabilistic guarantees that provide high-quality solutions. We perform extensive experiments showing that our methods outperform baselines. Furthermore, our algorithms scale on networks with up to billions of temporal edges, while baselines cannot handle such large networks. We use our techniques to analyze a financial network and show that our formulation reveals important network structures, such as bursty temporal events and communities of users with similar interests.
@@ 1. Introduction
Other promising applications captured by our formulation are, for example, (i) identifying frequent destinations appearing together from user travel data, i.e., groups of attractions visited following a specified travel journey captured by a temporal motif (Lei et al., 2020); and (ii) detecting dense groups of users and items according to specific purchase sequences e-commerce platforms, again captured trough temporal motifs (Pu et al., 2023; Cai et al., 2023). In both scenarios, the temporal motif-subnetwork densest subnetwork can provide unique insights, for (i) it can be used to design novel public transport routes or travel passes; and for (ii) it can be used for personalized advertising or identifying users with common purchase habits.
Solving the problem we propose requires exact enumeration of all temporal motif instances in the input network, which is extremely inefficient to perform. To address this issue we propose ALDENTE, a suite of randomized approximation ALgorithms for the temporal motif DENsest subneTwork problEm. ALDENTE contains two novel approximation algorithms leveraging a randomized sampling procedure, which has been successfully employed for temporal motif counting (Liu et al., 2019; Sarpe and Vandin, 2021b, a; Pu et al., 2023). Our novel algorithms can leverage any state-of-the-art unbiased sampling algorithm for temporal motif counting. The first algorithm in ALDENTE peels the vertex set in batches until it becomes empty (removing vertices participating in few temporal motif instances), estimating for small $\varepsilon>0$ within $(1\pm\varepsilon)$ the different counts needed for peeling. The second algorithm adopts a similar estimated batch-peeling technique but only for a fixed number of steps. On the remaining vertices it applies a greedy-peeling approach, where vertices that participate in few temporal motif instances are peeled one at a time. By carefully combining estimates, our algorithms output a high-quality solution with controlled error probability. Both our algorithms avoid exhaustive enumeration of all temporal motifs in the input network.
@@
Details on application scenarios for TMDS are provided in App. A, and a summary of the notation is found in App. B. We report in the Appendix, detailed proofs of our results (App. E), additional discussion on methods (App. D and App. F) and experiments (App. H). Our implementation is also available. [^1] https://github.com/iliesarpe/ALDENTE.
@@ 2. Preliminaries
(ii) a decaying function $\tau_{d}$: $\tau_{d}(S)=\frac{1}{\ell-1}\sum_{j=2}^{\ell}\exp(-\lambda(t_{j}-t_{j-1})),$ where $\lambda>0$ controls the time-exponential decay. In the above, $\tau_{c}$ is a simple weighting function where each $\delta$-instance receives a unitary weight, this can be used in most exploratory analyses where no prior information is known about the network. The decaying function $\tau_{d}$ is inspired from link-decay models, where the value of $\lambda$ corresponds to the inverse of the average inter-time distance of the timestamps in $E_{T}$. Therefore, the function $\tau_{d}$ accounts for decaying processes over networks, which are common in various domains, e.g., email communications, and spreading of diseases (Ahmed et al., 2021; Ahmad et al., 2021; Vazquez et al., 2007).
@@
We conclude with a fundamental definition, the temporal motif degree of vertex $v\in V_{T}$ is ${d}_{V_{T}}^{\tau}(v)\mathbin{=}\sum_{S\in\mathcal{S}_{V_{T}}}(\tau(S)\mathbbm{1}[v\in S]$), where $\mathbbm{1}[v\in S]$ is an indicator function denoting if the vertex appears at least once over the edges in the $\delta$-instance $S$. As an example, if we consider $T$ and $M$ as in Fig. 1a, $\tau_{c}$, and $\delta=10$, then ${d}_{V_{T}}^{\tau}(v_{4})=2$, as $v_{4}$ participates in two $\delta$-instances of $M$. We provide a notation table in Appendix B and all the missing proofs are deferred to Appendix E.
@@ 3. Related work
Densest-subgraph discovery in temporal networks. We do not discuss methods for DSD on temporal networks defined over snapshots of static graphs. Such a model is used for representing more coarse-grained temporal data than the fine-grained model we adopt, and it is less powerful (Holme and Saramäki, 2012).
@@ 4. Randomized algorithms
We now introduce the novel randomized approximation algorithms with tight probabilistic guarantees that we develop in ALDENTE. In fact, the algorithms in ALDENTE will address the TMDS avoiding the expensive exhaustive enumeration of the set $\mathcal{S}_{T}$, trough which the TMDS problem can be solved (see discussion in Section 5).
@@ 4.1. ProbPeel
The main idea behind ProbPeel is to avoid exhaustive enumeration of $\mathcal{S}_{T}$ by using random sampling and peeling vertices in “batches” according to an estimate of their temporal-motif degree. ProbPeel carefully combines highly-accurate probabilistic (1±$\pm\varepsilon,\varepsilon>0$) estimates of the temporal motif degrees of all vertices (that are obtained by suitably leveraging existing algorithms for global temporal motif estimation (Sarpe and Vandin, 2021b; Wang et al., 2020a)), and peels vertices in batches at each step of such a schema, similar non-approximate batch-peeling techniques have been employed successfully in other settings (Bahmani et al., 2012; Epasto et al., 2015; Sun et al., 2020a)).
@@
We assume access to a sampling algorithm $\mathcal{A}$, which given as input a subnetwork $T[W]$, where $W\subseteq V_{T}$, it outputs a sample $T^{\prime}=(V^{\prime},E^{\prime})\subseteq T[W]$. Given a subnetwork $T[W]\subseteq T$ we will execute the sampling algorithm $\mathcal{A}$ to obtain $r\in\mathbb{N}$ i.i.d. subnetworks $\mathcal{D}=\{{T}_{1},\dots,{T}_{r}:{T}_{i}=\mathcal{A}(T[W])\subseteq T[W],i\in[r]\}$. We can then use the subnetworks from $\mathcal{D}$ to compute an unbiased estimator $\hat{d}_{W}^{\tau}(w)$ of the temporal-motif degree for $w\in W$, for which, under the correct choice of $r$ it holds that “$\hat{d}_{W}^{\tau}(w)\in(1\pm\varepsilon){d}_{W}^{\tau}(w)$”, and combine such estimates to obtain an estimator $\hat{\tau}(W)\in(1\pm\varepsilon)\tau(W)$. Later, we will briefly mention how such estimators are computed. More details on the estimators are given in Appendix D.1.1.
@@ 8       $W_{i+1}\leftarrow W_{i}\setminus L(W_{i})$; $i\leftarrow i+1$;
return $W=\operatorname*{arg\,max}_{j=1,\dots,i}\{\hat{\tau}(W_{j})/|W_{j}|\}$.
@@
Bounding the sample size. Algorithm 1 can employ any unbiased sampling algorithm $\mathcal{A}:T\mapsto 2^{T}$ as subroutine (Liu et al., 2019; Wang et al., 2020b; Sarpe and Vandin, 2021b). We used PRESTO-A (Sarpe and Vandin, 2021b) for which the authors provide bounds on the number of samples $r$ for event “$\hat{\tau}(W_{i})\in(1\pm\varepsilon)\tau(W_{i})$” to hold with arbitrary probability $>1-\eta$ for $\eta\in(0,1)$. In our function $\mathtt{GetBound}$ we need a stronger result as we need $(1\pm\varepsilon)$-approximation for all temporal motif degree of vertices $w\in W_{i}$. Hence we bound individually, for each vertex, the probability that $\hat{d}_{W_{i}}^{\tau}(w)\notin(1\pm\varepsilon){d}_{W_{i}}^{\tau}(w)$ and combine such result with a union bound over all vertices (more details are in Appendix D.1.2).
@@
With probability at least $1-\eta$, ProbPeel achieves a $\frac{(1-\varepsilon)^{2}}{k(1+\xi)(1+\varepsilon)^{2}}$-approximation ratio for TMDS problem.
Note that if the parameter $\varepsilon$ is small (e.g., $\varepsilon\approx 10^{-3}$), then ProbPeel outputs a solution with an approximation ratio close to $\tfrac{1}{k(1+\xi)}$. In Appendix D.1.2 we discuss in details how to leverage existing state-of-the-art sampling algorithms, and their guarantees, to obtain a $(1\pm\varepsilon)$-approximation of the temporal motif degrees to be used by ProbPeel (in line 1), where we also report all the missing details of the functions GetBound and GetEstimates.
@@ 4.2. HybridPeel
[…]
@@
With probability at least $1-\eta$, HybridPeel is a $\frac{(1-\varepsilon)^{2}}{k(1+\xi)(1+\varepsilon)^{2}}$-approximation algorithm for TMDS problem when $J>1$, while for $J=1$ it achieves a $\frac{(1-\varepsilon)^{2}}{k(1+\xi)(1+\varepsilon)}$-approximation ratio.
Time complexity. The time complexity of the algorithm can be bounded by $\mathcal{O}({RT}_{\mathtt{ProbPeel}-J}(T)+{RT}_{\mathtt{Greedy}}(T[W_{J+1}]))$, i.e., the sum of the runtime ($RT$) of executing ProbPeel for $J$ iterations, and then the peeling procedure that we denote with Greedy on $T[W_{J+1}]$. The analysis follows from ProbPeel and our analysis of a baselines in Appendix F.2. By noting that after the $J$ iterations it holds that $|W_{J+1}|=o(n)$ the complexity is bounded by $\mathcal{O}\!(Jr_{\max}\hat{m}^{\ell}+k^{3}o(n)^{k}\log(o(n))).$ This highlights that first peeling the network with the randomized batch peeling technique enables HybridPeel to work on a significantly small network, over which to execute a more refined peeling algorithm, affording an exact enumeration of temporal motif instances on $T[W_{J+1}]$, leading to an overall scalable and practical algorithm.
@@
HybridPeel $\frac{(1-\varepsilon)^{2}}{k(1+\xi)(1+\varepsilon)^{2}}$ $\xi>0,\varepsilon,\eta\in(0,1),J>0$ $>1-\eta$ $\mathcal{O}\!\left(Jr(\varepsilon,\eta)\hat{m}^{\ell}+k^{3}o(n)^{k}\log(o(n))\right)$
@@ 6.2. Setup
Temporal motifs. The temporal motifs we use are good representative of a general input to TMDS in most applications, and are reported in Fig. 2. They represent different topologies, e.g., triangles, squares and more complicated patterns, spanning different values of $k$ and $\ell$, and the largest temporal motifs correspond to particularly hard inputs.
@@ 6.3. Solution quality and runtime
Concerning the solution quality we observe similar trends on most datasets. Batch provides solutions with smaller density than Greedy, and ProbPeel closely matches Batch’s results (as captured by our analysis). Greedy is the deterministic approximation algorithm that mostly often provides the solution with highest density when it terminates, and our HybridPeel closely matches such results, so our HybridPeel results accurate, scalable and efficient (as it is consistently faster than Greedy). Interestingly on the EquinixChicago dataset, where only our proposed randomized algorithms are able to terminate in less than three hours (ProbPeel takes few hundreds of seconds), the difference in the solutions provided by ProbPeel and HybridPeel are not significantly large, suggesting that on very large datasets ProbPeel may be a good candidate to obtain high-quality solutions in a very short amount of time.
@@ 6.4. Case study
The optimal solution $T[W^{*}_{\delta_{1}}]$ has 14 vertices, and we report its directed static network and its temporal support in Fig. 4 (Left). Interestingly, this is a star shaped network, with only the central vertex ($v_{1}$) exchanging money with all the other vertices (not reciprocated), and vertices are not friends. The message associated to the each transaction is identical for all transactions: “Sorry! We’re already sold out for tonight! Feel free to join us this even in the regular line and pay cover when you get there. Thanks!”. Even more interestingly, all events occur really close in time (see Fig. 4 (Left)). In fact, this corresponds to a bursty event with merchant $v_{1}$ overbooking for a specific event, which is identified by the combination of $\mathsf{M}_{5}$ and small $\delta\in\mathbb{R}^{+}$. We also observe that such a subnetwork cannot be captured by the existing formulations for dense temporal subnetworks (see Section 3) as both $T[W^{*}_{\delta_{1}}]$ and its directed static network have very small edge-density, i.e., $13/14<1$.

### v2

@@ Abstract.
In this work we introduce the novel problem of identifying the temporal motif densest subnetwork, i.e., the densest subnetwork with respect to temporal motifs, which are high-order patterns characterizing temporal networks. Identifying temporal motifs is an extremely challenging task, and thus, efficient methods are required. To address this challenge, we design two novel randomized approximation algorithms with rigorous probabilistic guarantees that provide high-quality solutions. We perform extensive experiments showing that our methods outperform baselines. Furthermore, our algorithms scale on networks with up to billions of temporal edges, while baselines cannot handle such large networks. We use our techniques to analyze a financial network and show that our formulation reveals important network structures, such as bursty temporal events and communities of users with similar interests.
@@ 1. Introduction
Other promising applications captured by our formulation are, for example, (i) identifying frequent destinations appearing together from user travel data, i.e., groups of attractions visited following a specified travel journey captured by a temporal motif (Lei et al., 2020); and (ii) detecting dense groups of users and items according to specific purchase sequences on e-commerce platforms, again captured trough temporal motifs (Pu et al., 2023; Cai et al., 2023). In both scenarios, the temporal motif-subnetwork densest subnetwork can provide unique insights, for (i) it can be used to design novel public transport routes or travel passes; and for (ii) it can be used for personalized advertising or identifying users with common purchase habits.
Solving the problem we propose requires exact enumeration of all temporal motif instances in the input network, which is extremely inefficient to perform. To address this issue we propose ALDENTE, a suite of randomized approximation ALgorithms for the temporal motif DENsest subneTwork problEm. ALDENTE contains two novel approximation algorithms leveraging a randomized sampling procedure, which has been successfully employed for temporal motif counting (Liu et al., 2019; Sarpe and Vandin, 2021b, a; Pu et al., 2023). Our novel algorithms can leverage any state-of-the-art unbiased sampling algorithm for temporal motif counting. The first algorithm in ALDENTE peels the vertex set in batches until it becomes empty (removing vertices participating in few temporal motif instances), estimating for small $\varepsilon>0$ within $(1\pm\varepsilon)$ the different counts needed for peeling. The second algorithm adopts a similar estimated batch-peeling technique but only for a fixed number of steps. On the remaining vertices it applies a greedy-peeling approach, where vertices that participate in few temporal motif instances are peeled one at a time. By carefully combining estimates, our methods output a high-quality solution with controlled error probability. Both our algorithms avoid exhaustive enumeration of all temporal motifs in the input network.
@@
Details on application scenarios for TMDS are provided in App. A, and a summary of the notation is found in App. B. We report in the Appendix, detailed proofs of our results (App. E), additional discussion on methods (App. D and App. F) and experiments (App. H). Our implementation is publicly available. [^1] https://github.com/iliesarpe/ALDENTE.
@@ 2. Preliminaries
(ii) a decaying function $\tau_{d}$: $\tau_{d}(S)=\frac{1}{\ell-1}\sum_{j=2}^{\ell}\exp(-\lambda(t_{j}-t_{j-1})),$ where $\lambda>0$ controls the time-exponential decay.
In the above, $\tau_{c}$ is a simple weighting function where each $\delta$-instance receives a unitary weight, this can be used in most exploratory analyses where no prior information is known about the network. The decaying function $\tau_{d}$ is inspired from link-decay models, where the value of $\lambda$ corresponds to the inverse of the average inter-time distance of the timestamps in $E_{T}$. Therefore, the function $\tau_{d}$ accounts for decaying processes over networks, which are common in various domains, e.g., email communications, and spreading of diseases (Ahmed et al., 2021; Ahmad et al., 2021; Vazquez et al., 2007).
@@
We conclude the preliminaries with defining the temporal motif degree of vertex $v\in V_{T}$ by ${d}_{V_{T}}^{\tau}(v)\mathbin{=}\sum_{S\in\mathcal{S}_{V_{T}}}(\tau(S)\mathbbm{1}[v\in S]$), where $\mathbbm{1}[v\in S]$ is an indicator function denoting if the vertex appears at least once over the edges in the $\delta$-instance $S$. As an example, if we consider $T$ and $M$ as in Fig. 1a, $\tau_{c}$, and $\delta=10$, then ${d}_{V_{T}}^{\tau}(v_{4})=2$, as $v_{4}$ participates in two $\delta$-instances of $M$. We provide a notation table in Appendix B and all the missing proofs are deferred to Appendix E.
@@ 3. Related work
Densest-subgraph discovery in temporal networks. We do not discuss methods for DSD on temporal networks defined over snapshots of static graphs. Such a model is used for representing more coarse-grained temporal data than the fine-grained model we adopt, and it is less powerful (Holme and Saramäki, 2012; Gionis et al., 2024).
@@ 4. Randomized algorithms
We now introduce the novel randomized approximation algorithms with tight probabilistic guarantees that we develop in ALDENTE. The algorithms in ALDENTE will address the TMDS avoiding the expensive exhaustive enumeration of the set $\mathcal{S}_{T}$, trough which the TMDS problem can be solved (see discussion in Section 5).
@@ 4.1. ProbPeel
The main idea behind ProbPeel is to avoid exhaustive enumeration of $\mathcal{S}_{T}$ by using random sampling and peeling vertices in “batches” according to an estimate of their temporal-motif degree. ProbPeel carefully combines highly-accurate probabilistic (1±$\pm\varepsilon,\varepsilon>0$) estimates of the temporal motif degrees of all vertices (that are obtained by suitably leveraging existing algorithms for global temporal motif estimation (Sarpe and Vandin, 2021b; Wang et al., 2020a)), and peels vertices in batches at each step of such a schema, similar non-approximate batch-peeling techniques have been employed successfully in other settings (Bahmani et al., 2012; Epasto et al., 2015; Sun et al., 2020a).
@@
We assume a sampling algorithm $\mathcal{A}$, which outputs a sample $T^{\prime}=(V^{\prime},E^{\prime})\subseteq T[W]$ of a subnetwork $T[W]$. Given a subnetwork $T[W]\subseteq T$ we will execute the sampling algorithm $\mathcal{A}$ to obtain $r\in\mathbb{N}$ i.i.d. subnetworks $\mathcal{D}=\{{T}_{1},\dots,{T}_{r}:{T}_{i}=\mathcal{A}(T[W])\subseteq T[W],i\in[r]\}$. We use the subnetworks in $\mathcal{D}$ to compute an unbiased estimator $\hat{d}_{W}^{\tau}(w)$ of the temporal-motif degree for $w\in W$, and for the correct choice of $r$ it holds $\hat{d}_{W}^{\tau}(w)\in(1\pm\varepsilon){d}_{W}^{\tau}(w)$, and combine these estimates to obtain an estimator $\hat{\tau}(W)\in(1\pm\varepsilon)\tau(W)$. More details on the estimators are given in Appendix D.1.1.
@@ 8       $W_{i+1}\leftarrow W_{i}\setminus L(W_{i})$; $i\leftarrow i+1$;
return $W=\operatorname*{arg\,max}_{j=1,\dots,i}\{\hat{\tau}(W_{j})/|W_{j}|\}$;
@@
Bounding the sample size. Algorithm 1 can employ any unbiased sampling algorithm $\mathcal{A}:T\mapsto 2^{T}$ as subroutine (Liu et al., 2019; Wang et al., 2020b; Sarpe and Vandin, 2021b). We used PRESTO-A (Sarpe and Vandin, 2021b) for which the authors provide bounds on the number of samples $r$ for event “$\hat{\tau}(W_{i})\in(1\pm\varepsilon)\tau(W_{i})$” to hold with arbitrary probability $>1-\eta$ for $\eta\in(0,1)$. In our function $\mathtt{GetBound}$ we need a stronger result as we need $(1\pm\varepsilon)$-approximation for all temporal motif degree of vertices $w\in W_{i}$. Hence we bound individually, for each vertex, the probability that $\hat{d}_{W_{i}}^{\tau}(w)\notin(1\pm\varepsilon){d}_{W_{i}}^{\tau}(w)$ and combine such result with a union bound over all vertices. (more details are in Appendix D.1.2).
@@
With probability at least $1-\eta$, ProbPeel achieves a $\frac{(1-\varepsilon)^{2}}{k(1+\xi)(1+\varepsilon)^{2}}$-approximation ratio for the TMDS problem.
Note that if the parameter $\varepsilon$ is small (e.g., $\varepsilon\!\approx\!10^{-3}$), then ProbPeel outputs a solution with approximation ratio close to $\tfrac{1}{k(1+\xi)}$. In Appendix D.1.2 we discuss in details how to leverage existing state-of-the-art sampling algorithms, and their guarantees, to obtain a $(1\pm\varepsilon)$-approximation of the temporal motif degrees to be used by ProbPeel (in line 1), where we also report all the missing details of the functions GetBound and GetEstimates.
@@ 4.2. HybridPeel
The design of HybridPeel is motivated by a drawback of ProbPeel: even for very accurate estimates of temporal motif degrees it may provide a sub-optimal solution to TMDS, since even for $\varepsilon$ close to 0, ProbPeel only converges to the approximation ratio of $\frac{1}{k(1+\xi)}$, while the peeling procedure we apply on $W_{J+1}$ provides denser solutions (close to $1/k$). The intuition behind HybridPeel is that, the first $J$ iterations are used to prune the network sufficiently, for the algorithm to maintain in $W_{J+1}$ a dense subnetwork $W^{\prime}\subseteq W_{J+1}$ attaining at least a $1/k$-approximation ratio. Therefore: (i) the algorithm is scalable, as it avoids the computation of $\mathcal{S}_{T}$, and (ii) is expected to output high-quality solutions.
@@
With probability at least $1-\eta$, HybridPeel is a $\frac{(1-\varepsilon)^{2}}{k(1+\xi)(1+\varepsilon)^{2}}$-approximation algorithm for the TMDS problem when $J>1$, while for $J=1$ it achieves a $\frac{(1-\varepsilon)^{2}}{k(1+\xi)(1+\varepsilon)}$-approximation ratio.
Time complexity. The time complexity of the algorithm can be bounded by $\mathcal{O}({RT}_{\mathtt{ProbPeel}-J}(T)+{RT}_{\mathtt{Greedy}}(T[W_{J+1}]))$, i.e., the sum of the runtime ($RT$) of executing ProbPeel for $J$ iterations, and then the peeling procedure that we denote with Greedy on $T[W_{J+1}]$. The analysis follows from ProbPeel and our analysis of a baselines in Appendix F.2. By noting that there exists a value $\gamma>1$ such that $|W_{J+1}|=\mathcal{O}(n^{1/\gamma})$ the complexity is bounded by $\mathcal{O}\!(Jr_{\max}\hat{m}^{\ell}+\gamma^{-1}k^{3}\,n^{k/\gamma}\log(n))$. This highlights that first peeling the network with the randomized batch peeling technique enables HybridPeel to work on a significantly small network, over which to execute a more refined peeling algorithm, affording an exact enumeration of temporal motif instances on $T[W_{J+1}]$, leading to an overall scalable and practical algorithm.
@@
HybridPeel $\frac{(1-\varepsilon)^{2}}{k(1+\xi)(1+\varepsilon)^{2}}$ $\xi>0,\varepsilon,\eta\in(0,1),J>0$ $>1-\eta$ $\mathcal{O}\!\left(Jr(\varepsilon,\eta)\hat{m}^{\ell}+\gamma^{-1}k^{3}\,n^{k/\gamma}\log(n)\right)$
@@ 6.2. Setup
Temporal motifs. The temporal motifs we use are good representative of a general input to TMDS in most applications, and are reported in Fig. 2. They represent different topologies, e.g., triangles, squares and more complicated patterns, spanning different values of $k$ and $\ell$, and the largest temporal motifs correspond to particularly challenging inputs.
@@ 6.3. Solution quality and runtime
Concerning the solution quality we observe similar trends on most datasets. Batch provides solutions with smaller density than Greedy, and ProbPeel closely matches Batch’s results (as captured by our analysis). Greedy is the deterministic approximation algorithm that mostly often provides the solution with highest density when it terminates, and our HybridPeel closely matches such results, so our HybridPeel results accurate, scalable and efficient (as it is consistently faster than Greedy). Interestingly on the EquinixChicago dataset, where only our proposed randomized algorithms are able to terminate in less than three hours (ProbPeel takes few hundreds of seconds), the difference in the quality of the solutions provided by ProbPeel and HybridPeel is not significantly large, suggesting that on massive data ProbPeel may be a good candidate to obtain high-quality solutions in a very short amount of time.
@@ 6.4. Case study
The optimal solution $T[W^{*}_{\delta_{1}}]$ has 14 vertices, and we report its directed static network and its temporal support in Fig. 4 (Left). Interestingly, this is a star shaped network, with only the central vertex ($v_{1}$) exchanging money with all the other vertices (not reciprocated), and vertices are not friends. The message associated to each transaction is identical for all transactions: “Sorry! We’re already sold out for tonight! Feel free to join us this even in the regular line and pay cover when you get there. Thanks!”. Even more interestingly, all events occur really close in time (see Fig. 4 (Left)). In fact, this corresponds to a bursty event with merchant $v_{1}$ overbooking for a specific event, which is identified by the combination of $\mathsf{M}_{5}$ and small $\delta\in\mathbb{R}^{+}$. We also observe that such a subnetwork cannot be captured by the existing formulations for dense temporal subnetworks (see Section 3) as both $T[W^{*}_{\delta_{1}}]$ and its directed static network have very small edge-density, i.e., $13/14<1$.